Two waves having the same frequency, wavelength, and amplitude are traveling in the same direction. If they differ in phase by π/2 and each has an amplitude of 0.060 m, what is the amplitude of the resultant wave?

cm
Answer #1

Let the waves be
   x_1=a\sin\theta ,~~x_2=a\sin(\theta +\phi)
where, \phi is the phase difference.
So, the resultant wave is
   x=x_1+x_2=a\sin\theta+a\sin(\theta +\phi)
\Rightarrow x=a\big[\sin\theta+\sin(\theta +\phi)\big]
  r-2α cos ( o/ 2) sin(θ+ o/2)
So, the amplitude of the resultant wave is given by
   a, = 2a cos(o/2)
So, given the values,
   6.0 cm, φ =-/2 a = 0.060 772
we get the resultant wave amplitude as
   a = 2 × 6.0 × cos(π/4) cm
  a-8.485 cm